# Models in Long-Context Interaction

### v1

@@ Appendix C Case Study
In this task, the model is given a long context in which an emotional segment is embedded within unrelated noise. The RAG method enables the model to retrieve a more accurate segment, leading to improved performance; CoEM further conducts emotional analysis on the retrieved segment, resulting in the greatest performance improvement.
@@
In this task, the model receives multiple emotional segments. The RAG method ranks the original segments based on their relevance, while CoEM further enhances the emotional features of the segments and ranks the enriched packs. This relevance-based ranking approach significantly boosts the model’s ability to distinguish emotions.
@@
In this task, we evaluate the model’s responses based on the F1 similarity with the ground truth. RAG helps the model retrieve more relevant source content, thereby improving its performance. However, the CoEM method, when introducing external knowledge, may alter certain internal details, which can lead to a drop in model performance. In Figure 12, we highlight the differences between the summary and the original text in red for comparison.
@@
In this task, the model is placed within a multi-turn dialogue context. The RAG method ranks the context chunks based on their relevance to the previous three dialogue turns. CoEM, after the initial ranking, generates a summary by combining the previous three turns with the initially selected chunks, and then performs a second round of relevance ranking between the initially filtered chunks and this summary, further ensuring the accuracy of the relevance assessment.
@@
In this task, the model is required to summarize specific characteristics of a psychological counseling report. RAG ranks the chunks based on their similarity to the target characteristics. CoEM further injects the analysis of these chunks provided by CoEM-Sage. However, since psychological counseling is a holistic process, analyzing only isolated chunks may lead to incorrect conclusions, resulting in a decline in model performance.
@@
In this task, the model is placed in an emotional situation, where it is required to answer the PANAS scale and express its emotions. RAG ranks the context chunks based on the query at each stage, while CoEM performs a finer-grained emotional analysis of these chunks. The CoEM-Sage model, with its stronger emotional intelligence (EI) capabilities, captures emotional cues more precisely, which in turn helps the tested CoEM-Core model better understand and express its own emotions.
Appendix D Synthetic Data Ablation
@@
Appendix E Details of RAG and CoEM
@@
We also report the chunk size and retrieved count for each task in Table 9. In QA, models use different chunk sizes. For EE, the retrieved counts correspond to stages 2–5.
Appendix F LLM as Judge Metrics Design
@@
In the Emotion Conversation task, we design metrics for each dialogue stage based on Cognitive Behavioral Therapy (CBT), Acceptance and Commitment Therapy (ACT), Humanistic Therapy, Existential Therapy, and Satir Family Therapy. The description and theoretical foundations for the design of each metric can be found in Table 10. Further explanations and practical applications of each metric can be referenced in the evaluation prompts, which are presented in Appendix H.
@@
Appendix G Unified Format of Data
We present data samples for each task in Figures 14 to 19. In the Emotion Classification task, the model analyzes the subject’s emotional state based on the given context. Emotion Detection requires the model to identify segments that carry distinct emotional expressions. In Emotion QA, the model answers questions grounded in contextual information. The Emotion Conversation task places the model in the role of a psychological counselor, responding to the client’s previous turn. Emotion Summary challenges the model to generate a structured summary of a counseling session, including the cause, symptoms, treatment process, illness characteristics, and treatment effect. Finally, in the Emotion Expression task, the model is immersed in an emotional situation, responds to the PANAS scale, and articulates its emotional state.
Appendix H Comprehensive Prompt Collections
This section presents the complete set of prompts used throughout the framework, encompassing Evaluation, Multi-agent Enrichment, and Emotional Ensemble Generation stages across all tasks. For tasks adopting automatic evaluation as the metric, we utilize GPT-4o as the evaluation model, with detailed evaluation prompts illustrated in Figures 20 to 25. During the Multi-Agent Enrichment stage, task-specific prompts are designed to guide agent collaboration and reasoning, as shown in Figures 26 to 31. Finally, in the Emotional Ensemble Generation stage, we employ carefully constructed prompts to support emotional diversity and coherence in response generation, with the full set depicted in Figures 32 to 37.
Figure 12: An overview of retrieved chunks across all tasks. In models’ final generation prompts, the Base setting includes none of the information; the RAG setting includes only the Content information; and the CoEM setting includes both the Content and Summary information.
Metric One-Stage Generation Two-Stage Generation
Establishing the Therapeutic Alliance 4.88 4.92
Emotional Acceptance and Exploration Guidance 4.36 4.38
Systematic Assessment 3.86 3.79
Recognizing Surface-Level Reaction Patterns 4.13 4.1
Deep Needs Exploration 4.13 4.32
Pattern Interconnection Analysis 3.66 3.77
Adaptive Cognitive Restructuring 3.60 3.73
Emotional Acceptance and Transformation 4.12 3.96
Value-Oriented Integration 3.94 3.69
Consolidating Change Outcomes and Growth Narrative 4.52 4.63
Meaning Integration and Future Guidance 4.16 4.19
Autonomy and Resource Internalization 4.84 4.86
Avg 4.18 4.20
Table 7: The comparison experiment results of synthetic data.
Figure 13: Dataset generation prompt for Emotion Conversation.
Task Chunking Initial Ranking Multi-Agent Enrichment Re-Ranking
EC Chunk by length Compute similarity with query GPT‑4o injects external EI analysis into each chunk Compute chunks’ similarity with query
ED Each segment as a chunk Skip this stage GPT‑4o injects external EI analysis into each chunk For each chunk, compute pairwise similarity with other chunks and select those with the lowest similarity
QA Chunk by length Compute similarity with query GPT‑4o injects external EI analysis into each chunk Compute similarity with the query
MC‑4 Chunk by length Compute similarity with query GPT‑4o generates an overall summary Compute chunks’ similarity with summary
ES Chunk by length Compute similarity with query GPT‑4o injects external EI analysis into each chunk Compute chunks’ similarity with query
$E$ Chunk by length Compute similarity with query GPT‑4o injects external EI analysis into each chunk Compute chunks’ similarity with query
Table 8: Application details in the CoEM framework.
Task Chunk Size Retrieved Count
EC 128 1
ED Num of segs 8
QA
GPT-4o-mini 128 8
GPT-4o 128 4
Deepseek-V3 512 4
Qwen3-8B 128 4
Llama-3.1-8B-Instruct 512 4
MC-4 128 4
ES 128 4
$E$ 128 2,4,4,4
(a) Parameter settings applied to RAG. For EE, the retrieved counts correspond to stages 2–5.
Task Chunk Size Count 1 Count 2
EC 128 1 1
ED Num of segs 8
QA
GPT-4o-mini 128 16 8
GPT-4o 128 16 4
Deepseek-V3 512 8 4
Qwen3-8B 128 16 4
Llama-3.1-8B-Instruct 512 8 4
MC-4 128 16 4
ES 128 10 4
$E$ 128 4,8,8,8 2,4,4,4
(b) Parameter settings applied to CoEM. Count 1 represents the number of chunks retrieved in the Initial Ranking stage, and Count 2 represents the number of chunks retrieved in the Re-Ranking stage.
Table 9: Parameter settings for RAG and CoEM approaches.
Stage Metric Name Description Theoretical Foundation Reception & Inquiry Establishing the Therapeutic Alliance Establish initial trust through empathy and a non-judgmental attitude, providing a safe foundation for further interventions. All theoretical orientations Emotional Acceptance and Exploration Guidance Guide the client to express emotions (e.g., anxiety, helplessness) in a safe atmosphere, demonstrating acceptance. Humanistic, ACT, Satir Family Systematic Assessment Integrate cognitive, behavioral, emotional, relational, and existential factors into a multidimensional assessment. CBT, Existential, Satir Family Diagnostic Recognizing Surface-Level Reaction Patterns Identify the client’s automatic cognitive, emotional, and behavioral responses. All theoretical orientations Deep Needs Exploration Reveal unmet psychological needs such as security, autonomy, connection, or meaning. All theoretical orientations Pattern Interconnection Analysis Understanding the interaction of problems within the individual’s internal systems and external systems; integrating findings from various dimensions to present a panoramic view of how the problem is maintained. All theoretical orientations Consultation Adaptive Cognitive Restructuring By examining the truthfulness and constructiveness of thoughts, build a more adaptive cognitive framework. CBT, Existential, Satir Family Emotional Acceptance and Transformation Developing Emotional Awareness, Acceptance, and Transformation Skills. All theoretical orientations Value-Oriented Integration Anchor Change to the Life Dimension Beyond Symptoms. All theoretical orientations Consolidation & Ending Consolidating Change and Growth Narrative Review therapeutic progress and reinforce positive change through a coherent personal narrative. All theoretical orientations Meaning Integration and Future Guidance Internalize therapy gains into a life philosophy and create a value-driven future plan. Humanistic, Existential, ACT Autonomy and Resource Internalization Strengthen the client’s internal coping resources and ability to continue growth independently. CBT, Existential, Satir Family
Table 10: Design of Emotion Conversation evaluation metrics. Theoretical Foundation denotes the source of inspiration for the design of this metric, and All theoretical orientations indicates that the metric draws upon all five theoretical orientations.

### v2

@@ B.2 Case Analysis of RAG and CoEM
In this task, the model is given a long context in which an emotional segment is embedded within unrelated noise. The RAG method enables the model to retrieve a more accurate segment, leading to improved performance; CoEM further conducts emotional analysis on the retrieved segment, resulting in the greatest performance improvement, as shown in Figure 13.
Figure 13: Case analysis of RAG and CoEM in Emotion Classification.
@@
In this task, the model receives multiple emotional segments. The RAG method ranks the original segments based on their relevance, while CoEM further enhances the emotional features of the segments and ranks the enriched packs. This relevance-based ranking approach significantly boosts the model’s ability to distinguish emotions. We skip the Initial-Ranking to capture richer emotional features. After enhancing the chunks with Multi-Agent Enrichment, we perform Re-Ranking to select the chunks that are least similar to others, as shown in Figure 14.
Figure 14: Case analysis of RAG and CoEM in Emotion Detection.
@@
In this task, we evaluate the model’s responses based on the F1 similarity with the ground truth. RAG helps the model retrieve more relevant source content, thereby improving its performance. Next, CoEM-Sage performs extraction on each retrieved chunk based on the query, retaining only the parts that are relevant to the query, as shown in Figure 15.
Figure 15: Case analysis of RAG and CoEM in Emotion QA.
@@
In this task, the model is placed within a multi-turn dialogue context. The RAG method ranks the context chunks based on their relevance to the previous three dialogue turns. CoEM, after the initial ranking, generates a summary by combining the previous three turns with the initially selected chunks, and then performs a second round of relevance ranking between the initially filtered chunks and this summary, further ensuring the accuracy of the relevance assessment, as shown in Figure 16.
Figure 16: Case analysis of RAG and CoEM in Emotion Conversation.
@@
In this task, the model is required to summarize specific characteristics of a psychological counseling report. RAG ranks the chunks based on their similarity to the target characteristics. CoEM further injects the analysis of these chunks provided by CoEM-Sage, as shown in Figure 17.
Figure 17: Case analysis of RAG and CoEM in Emotion Summary.
@@
In this task, the model is placed in an emotional situation, where it is required to answer the PANAS scale and express its emotions. RAG ranks the context chunks based on the query at each stage, while CoEM performs a finer-grained emotional analysis of these chunks. The CoEM-Sage model, with its stronger emotional intelligence (EI) capabilities, captures emotional cues more precisely, which in turn helps the tested CoEM-Core model better understand and express its own emotions, as shown in Figure 18.
Figure 18: Case analysis of RAG and CoEM in Emotion Expression.
B.3 Advantages of LongEmotion in Enhancing Emotional Intelligence
In this section, we discuss the advantages of the LongEmotion benchmark in promoting the full utilization of LLMs’ Emotion Intelligence capabilities in long-context interaction.
Psychological theories guided benchmark design.
In the Emotion Conversation task, we design scientifically rigorous evaluation metrics based on various psychological therapies and stages of dialogue data. For the Emotion Summary task, annotators summarize key elements of patient records considering physiological factors, personal growth history, and social factors, which can be seen in Table 5. In the Emotion Expression task, under given scenarios, models are guided to perform staged long-text self-expression in the rigorously designed framework.
Physiological Factors i) Biological, Genetic & Medical Factors. e.g., family medical history.
ii) Lifestyle Habits. e.g., sleep, diet, and exercise patterns.
Growth History i) Quality of interpersonal relationships during development.
ii) Academic and occupational performance during development.
Social Factors i) Family support system. e.g., emotional and financial support.
ii) Peer support system. e.g., friendship, social belonging and trust.
iii) Stressful life events. e.g., bereavement, job loss and daily stress.
Table 5: Annotation discipline for the annotation process of Emotion Summary.
Quality-guaranteed synthetic translation data.
@@ B.3 Advantages of LongEmotion in Enhancing Emotional Intelligence
Metric One-Stage Two-Stage
Establishing the Therapeutic Alliance 4.88 4.92
Emotional Acceptance and Exploration Guidance 4.36 4.38
Systematic Assessment 3.86 3.79
Recognizing Surface-Level Reaction Patterns 4.13 4.10
Deep Needs Exploration 4.13 4.32
Pattern Interconnection Analysis 3.66 3.77
Adaptive Cognitive Restructuring 3.60 3.73
Emotional Acceptance and Transformation 4.12 3.96
Value-Oriented Integration 3.94 3.69
Consolidating Change Outcomes and Growth Narrative 4.52 4.63
Meaning Integration and Future Guidance 4.16 4.19
Autonomy and Resource Internalization 4.84 4.86
Avg 4.18 4.20
Table 6: The comparison experiment results of synthetic data. One-Stage represents straightforward generation without the counseling note and the detailed skills. Two-Stage represents our generation method.
Comprehensive Experiments and In-Depth Case Studies.
We conducted extensive experiments on Base, RAG, and CoEM frameworks, accompanied by detailed case studies based on model outputs. Under the LongEmotion benchmark, various models exhibited distinct limitations—even the most advanced GPT-5 demonstrated issues such as overly mechanical responses despite its stronger theoretical capabilities.
Appendix C Details of RAG and CoEM
@@
Task Chunking Initial Ranking Multi-Agent Enrichment Re-Ranking
EC Chunk by length Compute chunk-query similarity External injection into each chunk Compute chunk-query similarity
ED Each segment as a chunk Skip this stage External injection into each chunk Select chunks with lowest similarity scores
QA Chunk by length Compute chunk-query similarity External injection into each chunk Compute chunk-query similarity
MC-4 Chunk by length Compute chunk-query similarity Generate an overall summary Compute chunk-query similarity
ES Chunk by length Compute chunk-query similarity External injection into each chunk Compute chunk-query similarity
$E$ Chunk by length Compute chunk-query similarity External injection into each chunk Compute chunk-query similarity
Table 7: Application details in the CoEM framework.
We also report the chunk size and retrieved count for each task in Table 8. In QA, models use different chunk sizes. For EE, the retrieved counts correspond to stages 2–5. The retrieved count of the one-time ranking in RAG is the same as the parameter settings for Re-Ranking in the table.
Task Chunk Size Initial Ranking Re-Ranking
EC 128 1 1
ED Num of segs 8
QA
GPT-4o-mini 128 16 8
GPT-4o 128 16 4
Deepseek-V3 512 8 4
Qwen3-8B 128 16 4
Llama-3.1-8B-Instruct 512 8 4
MC-4 128 16 4
ES 128 8 4
$E$ 128 4,8,8,8 2,4,4,4
Table 8: Parameter settings applied to CoEM. Initial Ranking and Re-Ranking denote the number of chunks retrieved in each respective stage.
Appendix D LLM as Judge Metrics Design
@@
In the Emotion Conversation task, we design metrics for each dialogue stage based on Cognitive Behavioral Therapy (CBT), Acceptance and Commitment Therapy (ACT), Humanistic Therapy, Existential Therapy, and Satir Family Therapy. The description and theoretical foundations for the design of each metric can be found in Table 10.
Stage Metric Name Description
Reception & Inquiry Establishing the Therapeutic Alliance Establish initial trust through empathy and a non-judgmental attitude, providing a safe foundation for further interventions.
Emotional Acceptance and Exploration Guidance Guide the client to express emotions (e.g., anxiety, helplessness) in a safe atmosphere, demonstrating acceptance.
Systematic Assessment Integrate cognitive, behavioral, emotional, relational, and existential factors into a multidimensional assessment.
Diagnostic Recognizing Surface-Level Reaction Patterns Identify the client’s automatic cognitive, emotional, and behavioral responses.
Deep Needs Exploration Reveal unmet psychological needs such as security, autonomy, connection, or meaning.
Pattern Interconnection Analysis Understanding the interaction of problems within the individual’s internal systems and external systems; integrating findings from various dimensions to present a panoramic view of how the problem is maintained.
Consultation Adaptive Cognitive Restructuring By examining the truthfulness and constructiveness of thoughts, build a more adaptive cognitive framework.
Emotional Acceptance and Transformation Developing Emotional Awareness, Acceptance, and Transformation Skills.
Value-Oriented Integration Anchor change to the life dimension beyond symptoms.
Consolidation & Ending Consolidating Change and Growth Narrative Review therapeutic progress and reinforce positive change through a coherent personal narrative.
Meaning Integration and Future Guidance Internalize therapy gains into a life philosophy and create a value-driven future plan.
Autonomy and Resource Internalization Strengthen the client’s internal coping resources and ability to continue growth independently.
Table 10: Design of Emotion Conversation evaluation metrics.
@@
Appendix E Unified Format of Data
We present data samples for each task in Figures 20 to 25. Emotion Detection requires the model to identify segments that carry distinct emotional expressions. In the Emotion Classification task, the model analyzes the subject’s emotional state based on the given context. In Emotion QA, the model answers questions grounded in contextual information. The Emotion Conversation task places the model in the role of a psychological counselor, responding to the client’s previous turn. Emotion Summary challenges the model to generate a structured summary of a counseling session, including the cause, symptoms, treatment process, illness characteristics, and treatment effect. Finally, in the Emotion Expression task, the model is immersed in an emotional situation, responds to the PANAS scale, and articulates its emotional state.
Appendix F Comprehensive Prompt Collections
This section presents the complete set of prompts used throughout the framework, encompassing Evaluation, Multi-agent Enrichment, and Emotional Ensemble Generation stages across all tasks. For tasks adopting automatic evaluation as the metric, we utilize GPT-4o as the evaluation model, with detailed evaluation prompts illustrated in Figures 26 to 31. During the Multi-Agent Enrichment stage, task-specific prompts are designed to guide agent collaboration and reasoning, as shown in Figures 32 to 37. Finally, in the Emotional Ensemble Generation stage, we employ carefully constructed prompts to support emotional diversity and coherence in response generation, with the full set depicted in Figures 38 to 43.
Figure 19: Dataset generation prompt for Emotion Conversation.
Figure 20: Emotion Detection dataset example.